# Mind the Gap Why Neural Memory Fails Under Semantic Density

### v1

@@ 7 Conclusion
This paper has characterized three fundamental problems with context-based AI memory that together explain why such systems can appear unreliable despite their impressive capabilities on other tasks.
The first problem concerns theoretical limits on storage capacity. The Stability Gap we documented shows that neural memory collapses to near-random accuracy as semantic density increases, while discrete memory maintains perfect retrieval regardless of corpus size. This gap arises from what we term the Orthogonality Constraint: when embedding keys share semantic overlap, interference during retrieval becomes mathematically inevitable in continuous vector space. Absent mechanisms that enforce pattern separation (such as discrete identity, explicit orthogonalization at write time, or other anti-cross-talk structure), overlap-driven interference is unavoidable in inner-product-addressed superposition memory.
The second problem concerns correctness failures in production systems. We found that production memory systems fail not from retrieval degradation—which is handled adequately at moderate scale—but from schema drift (inconsistent predicate naming during fact extraction) and version ambiguity (old values persisting alongside new ones after correction). Our multi-model comparison demonstrates that these failures affect all tested frontier models, including Claude Sonnet 4.5, GPT-5.2, and GPT-4o-mini, with schema drift persisting even within single sessions where models can observe their own prior extractions. This confirms that the failures reflect architectural limitations inherent to generative extraction rather than bugs in any particular model’s implementation.
@@
We introduced Knowledge Objects as a structured memory system addressing all three problems through complementary mechanisms: discrete hash-based storage prevents interference between facts, controlled vocabularies prevent schema drift during extraction, explicit version chains prevent ambiguity when facts are corrected, and selective retrieval achieves $O(1)$ cost per query regardless of corpus size.
Our findings establish that reliable, economical AI memory requires a hybrid architecture that divides labor according to computational strengths: neural weights for reasoning, pattern recognition, and generalization; discrete storage for precise factual recall; and a learned router to direct each query to the appropriate subsystem. The path to trustworthy AI memory runs not through ever-larger context windows or ever-denser weight matrices, but through principled separation of concerns that respects the fundamental capabilities and limitations of each storage mechanism.
@@
Several directions merit further investigation. First, scaling discrete memory to millions or billions of facts will require integration with approximate nearest neighbor search techniques that maintain sub-linear query time. Second, trust-weighted retrieval that accounts for source reliability, recency, and corroboration could improve accuracy on contested or time-sensitive facts. Third, tighter integration with production LLMs could reduce hallucination rates on factual queries by conditioning generation on retrieved Knowledge Objects. Finally, the relationship between the Orthogonality Constraint and other forms of interference in neural networks (such as catastrophic forgetting during continual learning) deserves theoretical investigation.
@@ A.2 KO Compiler Prompt
factual statements as structured triples.
@@ A.3 Operational Semantics
    Pass text to KO Compiler, receive list of triples

### v2

@@ 7 Conclusion
This paper identifies a fundamental geometric limit on neural memory and proposes the architecture that overcomes it.
The first contribution is theoretical: the Orthogonality Constraint establishes that online neural memory systems storing facts in shared continuous parameters will experience interference proportional to semantic density. This is not a limitation of current models that might be resolved through scaling, it is a mathematical property of inner-product retrieval in continuous space. We validate this constraint on real-world data across modalities: scientific measurements collapse to 0.02% accuracy at $\rho=0.96$, and images collapse to 0.05% at $\rho_{\text{intra}}=0.82$. The constraint has immediate architectural consequences: reliable episodic memory requires discrete addressing.
@@
The third contribution is architectural: Knowledge Objects provide the structured discrete storage that satisfies the Orthogonality Constraint. Hash-based addressing enables deterministic lookup and surgical updates. Controlled vocabularies prevent schema drift. Version chains ensure clean supersession. Production systems have adopted discrete storage but lack these structured semantics, they are halfway to the solution. Knowledge Objects complete the architecture.
The path forward is clear: structured storage (not unstructured text), controlled vocabularies (not free-form extraction), version chains (not append-only logs), and neural reasoning where appropriate. The Orthogonality Constraint is not a problem to be solved by bigger models or better training, it is a geometric fact that dictates architectural choices. Reliable AI memory requires a bicameral design: discrete symbolic storage for episodic facts, neural weights for semantic reasoning. This is not a limitation to be lamented but a clarification that enables building systems that actually work.
@@
Several directions merit further investigation. First, scaling discrete memory to millions or billions of facts will require integration with approximate nearest neighbor search techniques that maintain sub-linear query time. Second, trust-weighted retrieval that accounts for source reliability, recency, and corroboration could improve accuracy on contested or time-sensitive facts. Third, tighter integration with production LLMs could reduce hallucination rates on factual queries by conditioning generation on retrieved Knowledge Objects. Fourth, the biological analogy suggests a “sleep consolidation” mechanism: unlike context windows which are ephemeral, KOs provide a stable buffer that could enable periodic fine-tuning of the base model on frequently-accessed facts, gradually transferring episodic knowledge to parametric weights, the computational analog of hippocampal-to-cortical memory transfer during sleep. Finally, the relationship between the Orthogonality Constraint and other forms of interference in neural networks (such as catastrophic forgetting during continual learning) deserves theoretical investigation.
@@ A.2 KO Compiler Prompt
factual statements as structured facts.
@@ A.3 Operational Semantics
    Pass text to KO Compiler, receive list of facts
@@ A.4 Implementation Notes
Multi-Valued Predicates.
Some predicates are inherently multi-valued (e.g., a country has multiple neighbors). The simple hash scheme $\text{id}=\text{hash}(s\|p)$ stores only one value per (subject, predicate) pair. Production systems can address this through: (1) list-of-objects storage, where each key maps to a list rather than a single value; (2) composite keys that include a disambiguator, e.g., $\text{hash}(s\|p\|\text{ordinal})$; or (3) predicate-aware routing that uses embedding search for inherently multi-valued predicates while using hash lookup for functional predicates like “capital.”